The clarke family went sailing on a lake. Their boat averaged 6 kilometers per hour. The Rourke family took their outboard runabout for a trip on the lake for the same amount of time. their boat averaged 14 kilometers per hour. The rourke family traveled 20 kilometers farther than the clarke familt. How many hours did each family spend on their boat trip.
Answer by mananth(16946)

The clarke family went sailing on a lake. Their boat averaged 6 kilometers per hour. The Rourke family took their outboard runabout for a trip on the lake for the same amount of time. their boat averaged 14 kilometers per hour. The rourke family traveled 20 kilometers farther than the clarke familt. How many hours did each family spend on their boat trip.
T = hours
X = distance
6*T = X: Clarke family...........................1
14*T = X + 20: Rourke family......................2
...
Now plug the value of 6T into equation 2
14*T = X + 20
14T = 6T + 20
8T = 20 Divide both sides by 8
T = 2.5 hours
m.ananth@hotmail.ca